OpenStudy (anonymous):

Aaron, Sarah and Neil travel a distance of 54.25 miles on a road trip. Neil drives the car for the greatest distance. Of the remaining distance, Sarah drives 18.4 miles, while Aaron drives one-third of the remaining distance. Which equation can be used to determine the distance Neil drives, x? I have worked on this question for 40 minutes straight it's driving me nuts man...

OpenStudy (anonymous):

Neil's distance: x Sarah's distance: 18.4 Aaron: 1/3 [54.25 - (x + 18.4)]

OpenStudy (anonymous):

How did you find Aaron's distance? (Or equation)

OpenStudy (anonymous):

x + 18.4 + 1/3 [54.25 - (x + 18.4)] = 54.25 Make sense? Simplify the above and see what you get.

OpenStudy (anonymous):

"one-third of the remaining distance" 1/3 of the remaining disance 1/3 x [total distance - distance driven by Neil and Sarah" 1/3 [54.25 - (x + 18.4)]

OpenStudy (anonymous):

Oh! That makes much more sense!

OpenStudy (anonymous):

Lemme just simplify the other one

OpenStudy (anonymous):

Your question says "which equation". Is this multiple choice? Do you have the selections?

OpenStudy (anonymous):

Multiple Choice: 23(54.25−x)=18.4 (23)(54.25+x)=18.4 (23)(54.25)=18.4+x (32)(54.25−x)=18.4

OpenStudy (anonymous):

(The "23" is 2/3 sorry about that, also goes for 32, 3/2)

OpenStudy (anonymous):

I also have hints if that helps: Part 1 Try to find the expression for the distance that Aaron and Sarah drove between them. Total distance is 54.25 miles and Neil drove x miles. Then how much did Sarah and Aaron drive? Part 1 Of the remaining distance, Sarah drives 18.4 miles and Aaron a third of the remaining distance. What fraction of the distance did Sarah drive? Part 2 Two-third of (54.25−x) miles is 18.4. Express this sentence as an algebraic equation and solve for x.

OpenStudy (anonymous):

i'm sorry. i made a mistake. Aaron's distance should be 1/3 (54.25 - x) So that makes the equation: x + 18.4 + 1/3(54.25 - x) = 54.25

OpenStudy (anonymous):

it wouldn't make sense to say aaron drives 1/3 of remaining distance after Neil and Sarah. He drives 1/3 of the remaining distance just excluding Neil

OpenStudy (anonymous):

Thats absolutely fine! Dont worry about it! :D I really appreciate this so much!

OpenStudy (anonymous):

Also, could I combine the two x's? Or is there another step?